### v1

# {textblock}

Abstract
Background. The recent surge in phishing attacks keeps undermining the effectiveness of the traditional anti-phishing blacklist approaches. On-device anti-phishing solutions are gaining popularity as they offer faster phishing detection locally.
Aim. We aim to eliminate the delay in recognizing and recording phishing campaigns in databases via on-device solutions that identify phishing sites immediately when encountered by the user rather than waiting for a web crawler’s scan to finish. Additionally, utilizing operating system-specific resources and frameworks, we aim to minimize the impact on system performance and depend on local processing to protect user privacy.
Method. We propose a phishing detection solution that uses a combination of computer vision and on-device machine learning models to analyze websites in real time. Our reference-based approach analyzes the visual content of webpages, identifying phishing attempts through layout analysis, credential input areas detection, and brand impersonation criteria combination.
Results. Our case study shows it’s feasible to perform background processing on-device continuously, for the case of the web browser requiring the resource use of 16% of a single CPU core and less than 84MB of RAM on Apple M1 while maintaining the accuracy of brand logo detection at 46.6% (comparable with baselines), and of Credential Requiring Page detection at 98.1% (improving the baseline by 3.1%), within the test dataset.
Conclusions.  Our results demonstrate the potential of on-device, real-time phishing detection systems to enhance cybersecurity defensive technologies and extend the scope of phishing detection to more similar regions of interest, e.g., email clients and messenger windows.
Index Terms:
Phishing, Computer Vision, macOS
1 Introduction
According to the phishing activity trends reports [^1] https://apwg.org/trendsreports/ from the Anti-Phishing Working Group (APWG), the number of phishing attacks keeps increasing year over year, with 2023 being the worst so far, with 4,987,809 unique phishing webpages created. While cloud infrastructure providers like Cloudflare implement phishing detection on their end [^2] https://blog.cloudflare.com/2023-phishing-report, these measures are not sufficient to prevent phishing attacks at large, so there is still a need for on-device solutions, both for desktop and mobile.
Existing phishing detection methods can be categorized into three categories: blacklist-based, classification-based, and reference-based. Traditional blacklist-based solutions like Google Safe Browsing  [^3] https://safebrowsing.google.com, although they are effective, do not keep up with the speed of phishing websites spreading [19], whose creation and deployment are typically automated. Classification-based approaches [18, 17, 10] use machine learning algorithms to analyze URL, HTML, or other features to classify webpages as phishing or legitimate, including on-device approaches [7, 2], but lose efficiency against HTML obfuscation techniques and lack accuracy. In contrast, reference-based approaches [1, 13, 14] use computer vision and can effectively analyze webpage appearance to extract information about the displayed content, allowing to detect phishing attempts and form reasonable verdicts.
While the latest reference-based solutions effectively detect phishing across the web, they focus on automated web crawling and filling the blacklists. The time after a phishing webpage is published on the web and remains unprocessed by detection systems is critical. A phishing campaign starts spreading malicious URLs just when they become available, and even within a few minutes, the campaign may affect many users.
In this paper, we propose a real-time on-device unobtrusive anti-phishing solution for macOS and evaluate in on the case of in-browser phishing website detection, improving upon the work of [14]. Our approach relies on macOS-specific system resources and frameworks, which allows us to significantly reduce computational resource demand. Processing live screen capturing requires applying a phishing detection algorithm within a matter of seconds, ensuring users’ safety. To achieve this, we rely solely on local machine learning models, eliminating the need for cloud-based tools. Another important aspect of using only local models is privacy concerns. By utilizing local models, user data remains on the device at all times, ensuring enhanced security and peace of mind for the user, which aligns with Apple’s approach to privacy protection [^4] https://www.apple.com/ua/privacy/approach-to-privacy/.
2 Threat Model
Figure 1: General workflow of our solution
We consider the threat model of PhishIntention [14], in which the adversary creates a phishing website, impersonating an official website of a known company and demanding the personal data of the website visitor. This form of phishing is executed through social interaction, employing psychological manipulations to trick users into disclosing sensitive security data. Initially, the attacker conducts a thorough analysis to identify the potential vulnerabilities within the target demographic essential for the successful execution of the attack. Subsequently, the phisher endeavors to establish trust with the target. In the final phase, the attacker manipulates the scenario to induce the target to share critical information.
3 Solution
Figure 1 presents a high-level overview of our proposed approach. Similarly to PhishIntention, we focus on identifying both the intention to capture credentials and the intention to impersonate a brand. We begin with the webpage layout analysis (Section 3.1). In the next step, we classify the webpage elements into two categories: credential required layout or not (Section 3.2). The final step is to classify the brand that the phishing webpage is trying to impersonate well-known brand (Section 3.3).
Data collected in these steps will help us to verify all three conditions: (a) verifying that the webpage layout is designed to collect credentials, (b) confirming that a detected brand appears on our reference list, and (c) ensuring that the webpage does not belong to one of the official brand domains. Matching on all three conditions allows us to conclude that the current page is attempting to phish users.
3.1 Abstract Layout Detection
The main purpose of abstract layout detection is to determine whether the currently visible webpage contains credential input forms that may be targeted by hackers and find the brand logo bounding box. To achieve this, we analyze screenshots, identifying different layout elements such as logos, buttons, inputs, labels, and blocks.
We have used the part of the dataset from the section CRP classifier and AWL detector of [15]. The dataset contains about 9K webpage screenshots, annotated with regions and types of layout elements, together with the information about the webpage layout type (credential required or not) and the corresponding brand.
Layout Elements Extraction
The DETR (Detection Transformer) [5] architecture revolutionizes object detection by employing transformers, to directly learn the relationships between objects in an image without predefined anchor boxes. Considering its state-of-the-art performance in object detection tasks, we have decided to adopt the DETR model with ResNet-50 [8] backbone. We trained the model using 8,109 labeled webpage screenshots. During the training stage, we conducted a comparative analysis of the DETR model’s performance across varying image sizes. Table I shows the mAP (Mean Average Precision) accuracy metric compared with processing speed. The model which works with small images shows good performance, albeit with lower accuracy.
TABLE I: DETR Performance Comparison
Image Size Logo mAP [.5:.95] Samples per Second
224x386 37.3 10.5
432x768 46.6 3.9
To safeguard user security, it’s imperative that our system processes webpage layout analysis at a minimum rate of once per second. Concurrently, it’s essential to manage the workload on the CPU effectively to prevent overutilization, ensuring the system remains responsive and reliable over extended periods of operation. Achieving an average processing rate of 3.9 samples per second provides the capability to analyze a frame in approximately 1/4 seconds. This rapid processing affords us a substantial interval of 3/4 seconds where the CPU can enter a reduced activity state, thereby mitigating the risk of overheating and sustaining performance over time.
Reducing Overlapping Boxes
The NMS (Non-Maximum Suppression) algorithm is used in object detection to eliminate redundant or overlapping bounding boxes. By design, the DETR model doesn’t have a built-in NMS layer. However, given that an abstract webpage layout should never contain overlapping elements, we integrated the protobuf model from coremltools  [^5] https://apple.github.io/coremltools/v3.4/index.html and merged it with the main model. DETR consistently generates a fixed set of 100 bounding boxes for any input image, which is the default value but can be changed. These bounding boxes are a representation of potential object locations within the image (Figure 2).
Figure 2: Potential object locations
NMS systematically identifies and keeps only the non-overlapping bounding boxes with the highest confidence scores and makes our outputs flexible, meaning now we can receive from 0 to 100 boxes instead of a fixed count. The remaining bounding boxes represent the most likely object locations within the image (Figure 3).
Figure 3: Refined object locations after post-processing
3.2 Credential Required Page Classification
When we have successfully extracted the layout from the webpage image, we also need to classify this layout into credential or non-credential required classes. To streamline our process and optimize algorithm we aimed to utilize already extracted features from Abstract Layout Detection step. The DETR model receives the query image as an input and generates N output vectors. Because output vectors are used to predict the class and location of the object, they consist of the information about the object and are suitable for the object-level image representation [4].
The object detection model is now divided into two models, the first being LayoutFeatureExtractor, which generates object-level features of shape (B, N, H) where B is batch size, N is a number of output vectors which is 100 in our case and H is hidden states for each output which is 256 in our case. The subsequent LayoutObjectDetector model takes these raw features as input and transforms it into a tensor of shape (B, N, 5) where the first 4 elements are coordinates of the bounding boxes and the last element is a layout element class. Then, we employ an additional LayoutClassifier model, which repurposes the previously extracted object-level features to transform the task of object detection into image classification. This transition is facilitated by a custom classification head (Figure 4), comprising a streamlined multi-layer perceptron architecture. The classification head processes and flattens the object-level features, applies ReLU activation functions for non-linear processing, and concludes with a Dropout layer to mitigate overfitting. Resulting in the output of shape (B, 2), this classification head efficiently provides classification scores, distinctly categorizing pages as credential required or not.
Figure 4: Classification head for DETR object-level features
With frozen weights of LayoutFeatureExtractor, we exclusively trained the classification head to transform the extracted features into classification scores. This approach allowed us to save disk storage and reduce RAM consumption by reusing already extracted features and has also greatly improved accuracy (Section 4.1.2).
3.3 Logo Brand Classification
We use brand logos available in the CRP classifier and AWL detector dataset part of [15]. From the available brands in the dataset, we select 15 brands known worldwide. We crop the logo elements from the entire screenshot and associate them with the corresponding brands. We used the ResNet-18 model, which has been trained on precisely cropped logo layout elements during the training stage. During the evaluation stage, we utilize the layout elements identified by the LayoutObjectDetector (DETR) and select the logo element with the highest confidence score. Subsequently, the LogoBrandClassifier (ResNet-18) is applied within the bounds of the detected logo’s bounding box. Accuracy metrics details are available in the Section 4.1.3.
3.3.1 Phishing detection
We used Netify [^6] https://www.netlify.com/ to get primary domains for our detectable brands. This list doesn’t contain all the brands that we’re detecting, and some had to be collected manually. Engaging in this process provided profound insights into the mechanisms through which phishing operations can be successfully executed. For example, DHL does not provide a publicly accessible set of its official domains. This lack of transparency is compounded by the existence of multiple domains, such as dhl.com, express.dhl, mydhli.com, dhlsameday.com, and dhlexpresscommerce.com, which could potentially confuse users and expose them to phishing risks.
3.4 Models Hierarchy
The resulting hierarchy of the final models is illustrated in Figure 5, showing the structured organization and interrelations among the different models. A, B and C sections represents different training stages. A is a initial stage where we are training LayoutFeatureExtractor simultaneously with LayoutObjectDetector on object detection task. B and C can be trained in parallel, but only when the previous stage was fully trained. When we make changes in the A stage, then stage C, which includes LayoutClassifier, should be retrained because it fully depends on LayoutFeatureExtractor outputs.
Figure 5: Final Models Hierarchy
3.5 macOS Application
All models were converted to the Core ML format to optimize performance through the use of Metal Performance Shaders. Core ML  [^7] https://docs.developer.apple.com/documentation/coreml is a built-in high-level macOS native framework to integrate machine learning models into Apple platform applications. It is optimized to perform on-device operations by leveraging the CPU, GPU, and Neural Engine chip while minimizing its memory footprint and power consumption. Given the need for models to adhere to the proprietary format (.mlmodel) for use, Core ML accommodates conversion from third-party frameworks through Model Intermediate Language  [^8] https://apple.github.io/coremltools/docs-guides/source/model-intermediate-language.html (MIL). Metal  [^9] https://docs.developer.apple.com/documentation/metal/ framework allows applications to directly interact with a device’s GPU. Machine learning applications exploit Metal’s computational acceleration for both training and inference tasks. Metal Performance Shaders  [^10] https://docs.developer.apple.com/documentation/metalperformanceshaders/ is a suite of highly optimized compute and graphics shaders engineered for seamless and efficient integration within Metal applications. These shaders are data-parallel primitives that are fine-tuned to exploit the unique hardware capabilities of each GPU family, maximizing performance.
We built a native macOS application using Swift programming language and integrated all converted Core ML packages into it. We utilized Screen Capture Kit to access fullscreen frames. Screen Capture Kit  [^11] https://docs.developer.apple.com/documentation/screencapturekit is a framework for screen capturing on macOS. This framework has been designed with a focus on performance by GPU utilization. It allows capturing content from various sources such as displays, applications, and windows, along with associated audio. A native macOS application continuously runs in the background, waiting for the active web browser to appear on the screen. While working with different browsers, we need to specify the region of actual web content and ignore the desktop background, address input, or tabs. The Accessibility Framework  [^12] https://docs.developer.apple.com/documentation/accessibility/ incorporates an extensive array of tools and functionalities aimed at enhancing the accessibility of Apple devices. For developers, this framework offers the opportunity to delve into the accessibility graph that contains all UI elements and their relations, providing an additional source of information regarding the active application’s state, layout, and controls. In our case, we can inspect the accessibility elements graph of Safari and find an element called AXWebArea, which contains the actual webpage. By getting its parent, which is AXScrollArea, we can account for the relative coordinates of the web area to the inner page scroll offset. These coordinates are used to crop the fullscreen image and resize it to 432x768 size, making the image acceptable for the machine learning models analysis. After this, the algorithm works exactly as described in Section 3.
4 Evaluation
We perform a case study with the Safari web browser, exploring several use aspects to ensure that our system is optimized for both performance and efficiency:
Abstract Layout Detection Efficiency. How does our model’s logo-detecting efficiency compare to the previous solutions in terms of both accuracy and processing speed?
Credential Required Page Classification Efficiency. What is the effectiveness of our object-level features-based classification compared to previous work?
Logo Brand Classification Efficiency. How accurately can we classify the brand of detected logo elements as part of a phishing detection system?
System Resources Usage. Can we perform continuous processing in the background without a major impact on user workflow?
4.1 Accuracy Metrics
4.1.1 Abstract Layout Detection
One of the main goals of layout detection is to find logo elements, so we have checked how well DETR does this task. The mAP (Mean Average Precision) measurement is under the IoU (Intersection over Union) thresholds [0.5:0.95] and shows slightly worse accuracy compared to the PhishIntention model, but at the same time, achieves the same results as Phishpedia [13] model (Table II). We achieved worse results than competitors, but our solution is more efficient in terms of processing speed.
TABLE II: Logo detection accuracy comparison
Solution Logo mAP [.5:.95] Samples per Second
PhishIntention 59.5 0.7
Phishpedia 46.6 -
Ours 46.6 3.9
4.1.2 Credential Required Page Classification
To validate the accuracy and robustness of object-level features based classification, we conducted a comparative analysis against the PhishIntention baseline (Table III). The higher test accuracy of our solution (98.1%) underscores its effectiveness in accurately classifying webpages, reflecting its robustness against overfitting and its enhanced ability to generalize from the training data to real-world applications. This increment (+3.1%) in test accuracy is particularly noteworthy as it indicates a significant reduction in false positives and false negatives, which are critical in the context of cybersecurity solutions.
TABLE III: CRP Classification Accuracy Comparison
Solution Train Accuracy Test Accuracy
PhishIntention 99.3 95.0
Ours 99.5 98.1
4.1.3 Logo Brand Classification
Determining a phishing website’s brand affiliation significantly depends on the accurate recognition and extraction of logos. This crucial phase of brand logo identification is substantially influenced by the performance of the underlying object detection model, particularly in accurately locating the logo’s bounding box. Table IV presents the accuracy metrics achieved by employing the ResNet-18 model for logo brand classification. These metrics illustrate the model’s performance in terms of its F1 score, precision, and recall. The combination of DETR for object detection and ResNet-18 model for image classification demonstrates good performance with an F1 score of 0.95, indicating a high degree of accuracy in the balanced measurement of precision and recall.
TABLE IV: Logo Brand Classification Accuracy
Solution F1 Precision Recall
DETR + ResNet-18 0.95 0.92 0.95
4.2 System Resources Usage
We measured performance on two different devices: MacBook Pro 2020 M1 and 4K iMac 2020 Intel Core i7 3.8 GHz. Initial tests were performed without imposing a frame rate limit, enabling the application to analyze each frame as it becomes available. The results shown in Table V highlight the performance difference between Apple’s M1 and Intel’s Core i7 processors, with the former demonstrating better optimization in terms of both CPU usage and RAM consumption.
TABLE V: System Resources Usage
Device CPU Usage RAM Consumption
Apple M1 58% 440 MB
Intel Core i7 74% 860 MB
To meet the criteria for unobtrusive background processing, we restrict the frame rate to a single frame per second. This measure is posited to strike an optimal balance between enhancing user security and minimizing the consumption of system resources. We observe (Table VI) a significant improvement in CPU utilization attributed to increased periods of processor idleness.
TABLE VI: System Resources Usage (FPS Lock)
Device CPU Usage RAM Consumption
Apple M1 20% 440 MB
Intel Core i7 11% 820 MB
Quantization [^13] https://huggingface.co/docs/optimum/concept_guides/quantization is a process of clamping the model weights. We quantized all models from a default 32-bit floating point to a smaller but less accurate 16-bit floating point. This optimization technique effectively halves the model’s size while maintaining a minimal impact on its accuracy. Using model quantization (Table VII) yields substantial enhancements in frame processing time, CPU utilization, and, most notably, RAM consumption. Furthermore, this process significantly reduces our app bundle size, decreasing from 254 MB to 136 MB.
TABLE VII: System Resources Usage (Quantized Models)
Device CPU Usage RAM Consumption
Apple M1 16% 84M
Intel Core i7 8% 450M
Our final metrics of system resource usage indicate that our solution can operate continuously in the background without a noticeable impact. It is important to highlight that CPU usage represents a fraction of the total available capacity: given that the Apple M1 has 8 cores, this means we are utilizing 16% of the 800% total capacity available across all cores. In comparison to other applications, this level of resource consumption is equivalent to maintaining three active Safari browser tabs (which typically consume about 5% depending on the website) or conducting an active Zoom call (which can use about 22% during a voice call and about 41% during video call).
5 Related Work
Reference-based phishing webpage detection
State-of-the-art approaches, such as PhishIntention [14], Phishpedia [13], and VisualPhishNet [1], primarily focus on automatically identifying phishing websites on the internet using web crawlers rather than relying on local inference. Solutions, such as DynaPhish [16] and KnowPhish Detector [12], further improve accuracy and extend the applicability of reference-based approaches but still are cloud-first. Our work addresses the performance aspect of such approaches, bringing them locally to the user devices without compromising on phishing detection accuracy.
In the original PhishIntention algorithm, Faster R-CNN [20] model was chosen, but we encountered different errors during attempts to convert this architecture to Core ML format. We have initially added support of a few operators to MIL, but after deeper investigation, it became clear that Faster R-CNN isn’t convertible to Core ML format because it requires a number of custom layers that are not available in Metal and Core ML, so we decided to continue with DETR architecture.
On-device phishing detection
Both industry practitioners and academia researchers have recently begun to explore this attack vector more actively, though, to the best of our knowledge, the current industry focus is on SMS phishing (smishing) prevention. Recent works from Samsung R&D Institute focus on smishing: Harichandana et al. [3] introduce an on-device pipeline for smishing detection, and J. W. Seo et al. [21] present a privacy-preserving SMS classifier. The privacy-preserving aspect of on-device solutions is also explored by researchers in recent works on SMS spam and phishing detection [22, 6].
6 Discussion
Proactive Protection
Our solution aims to automatically detect phishing attacks, eliminating the need for additional user interaction. There’s no need for users to press extra buttons or copy and paste suspicious URLs for detection. In stressful situations, users may lack the necessary attention to analyze URLs and page content thoroughly. Hence, our solution operates proactively, safeguarding users without their active intervention.
On-device Processing
One of the primary benefits of on-device processing is its ability to provide the fastest response to previously unknown threats without waiting for database updates. Our proposed solution can continuously perform processing in the background without extensive system resource usage. However, different hardware provides different levels of performance due to specific optimizations. An interesting finding was that the Intel processor demonstrates worse results in terms of CPU usage when the frame rate is not limited (Table V), but it becomes twice as efficient when the frame limit is set to the one (Table VI). This finding suggests we need to look more into how hardware acceleration works and test more devices to understand this behavior better.
Browser Extensions
In our approach, we have access to the image of the whole screen but select a region of interest, which we currently limit to an active browser, performing any screen recognition tasks within this window. In the use case of browser window content tracking, this approach has a direct advantage over relying on browser extensions, as it is browser-agnostic and requires minimal adjustment for working with different browsers: adding Google Chrome support on the Safari-compatible base required up to 50 lines of code only. For the users, the browser-agnostic approach has the direct advantage of lessening the burden of browser extension management.
Users as Sensors
Our solution combines with the traditional method of keeping a blacklist of phishing websites. The algorithm can be extended to perform URL matching with a blacklist at the very beginning. If there is no match in the blacklist, we can perform the main part of the algorithm described in this paper. Upon identifying a phishing attempt, both the user and our centralized system are alerted. The primary advantage of this method is that it leverages thousands of users as sensors rather than relying on multiple web crawlers to search for phishing attempts on the web, aligning with the human-as-a-security-sensor (HaaSS) [9] approach.
7 Limitations
The main limitations of our proposed solution stem from the specifics of the use of computer vision techniques for object detection on the webpage. Our approach only identifies phishing websites that display both a visible brand logo and a credential form on the same page. If either of these elements is missing or becomes hidden when a user scrolls or magnifies, our solution will fail.
We only used a small subset of well-known brands for our research, which means our solution doesn’t protect users from phishing attacks for unknown brands. Such an approach also fails when the company rebrands. For example, from Facebook to Meta or from Twitter to X. In case of the need for a well-known brand logos database update, machine learning models should be retrained and updated on the user device. A recent study [11] has also shown a possible attack on techniques that rely on brand logo recognition, for which we are not robust.
8 Conclusions and Future Work
This work argues for the potential of on-device, real-time phishing detection systems to enhance cybersecurity defensive technologies. We demonstrate that modern devices are capable of leveraging machine learning models efficiently, sustaining fast processing speed while conserving system resources. A straightforward future work direction is adapting our solution for iOS devices. This transition is expected to be relatively smooth, as the Core ML packages required for the transfer are fully compatible with the iOS platform.
In our case study, we focused on webpage phishing detection. However, given our capability to capture the entire screen, our potential for advancement is not limited here. A complex solution could orchestrate various algorithms tailored to specific Regions of Interest. For example, in web browsers, we can process webpages according to the introduced algorithm, in email clients, we can detect phishing emails, in messengers, we can detect smishing.
9 Acknowledgements
We would like to thank the anonymous reviewers for providing valuable feedback and suggestions that helped to improve this paper.